# Mitigating Hallucinations with Knowledge Editing

### v1

@@ 1 Introduction
Large Language Models (LLMs) have shown superior performance in various tasks (Zhao et al., 2023). However, one critical weakness is that they may output hallucinations, referring to the non-factual information in generated content, for reasons such as the limit of models’ internal knowledge scope or fast-changing world facts (Zhang et al., 2023). Considering the high cost of retraining LLMs from scratch, knowledge editing has been designed as a new paradigm to correct erroneous or outdated factual knowledge in LLMs (Wang et al., 2023b).
@@
Although there are many existing question-answering datasets such as $\text{WikiData}_{\text{recent}}$ (Cohen et al., 2024), ZsRE (Yao et al., 2023), and WikiBio (Hartvigsen et al., 2024) widely used for knowledge editing evaluation, one common issue is that they do not verify whether LLMs, before applying knowledge editing, actually generate hallucinated answers to the evaluation questions. When such datasets are adopted to evaluate the performance of LLMs after being edited, it is hard to directly use the scores to judge the effectiveness of different knowledge editing techniques in correcting hallucinations, which is the motivation of applying knowledge editing to LLMs. To better illustrate this point, following the evaluation setting in  (Zhang et al., 2024b), we conducted a preliminary study to examine the pre-edit and post-edit performances of Llama2-7B on the aforementioned three evaluation datasets. As shown in Table 1, we can clearly observe that Llama2-7B achieves a relatively high performance, measured by the rate of answering the evaluation questions correctly (Accuracy (%)), even before applying knowledge editing techniques. Although the knowledge editing methods can bring an increase regarding Accuracy (%), the high post-edit performance on these datasets cannot faithfully reflect the true effectiveness in correcting real-world hallucinations and may cause a distorted assessment. Thus, the fundamental question remains insufficiently validated: Can knowledge editing really correct hallucinations in LLMs?
@@
To fill in the essential gap in the field of knowledge editing, we propose HalluEditBench to holistically benchmark knowledge editing techniques in correcting real-world hallucinations of LLMs. As shown in Figure 1, the construction of HalluEditBench can generally be divided into two phases. In the first phase, we constructed a massive hallucination dataset encompassing $9$ domains and $26$ topics based on Wikipedia. For each of Llama2-7B, Llama3-8B, and Mistral-v0.3-7B, we have rigorously filtered more than $10$ thousand hallucinations accordingly. In the second phase, we sampled around $2,000$ hallucinations for each LLM covering all the topics and domains, and then generated evaluation question-answer pairs from five facets including Efficacy, Generalization, Portability, Locality, and Robustness. Through extensive empirical investigation on performance of $7$ typical knowledge editing techniques, including FT-L (Meng et al., 2022), FT-M (Zhang et al., 2024b), MEMIT (Meng et al., 2023), ROME (Meng et al., 2022), LoRA (Hu et al., 2022), ICE (Zheng et al., 2023), and GRACE (Hartvigsen et al., 2024), regarding the aforementioned five dimensions, we have provided novel insights into their potentials and limitations. A summary of the insights is as follows:
@@
  • No editing methods can outperform others across five facets and the performance beyond Efficacy for all methods is generally unsatisfactory. Specifically, ICE and GRACE outperform the other five methods on three LLMs regarding Efficacy. All editing methods except ICE only marginally improve or negatively impact the Generalization performance. Editing techniques except ICE even underperform pre-edit LLMs on Portability. FT-M and ICE surpass others on Locality performance. ICE has a poor Robustness performance compared to other methods.
@@ 2.2 Evaluation QA Pair Generation and Metrics
After constructing the hallucination dataset, we proposed to holistically assess the performance of knowledge editing methods in correcting hallucinations from five facets including Efficacy, Generalization, Portability, Locality, and Robustness. First, we leveraged GPT-4o to generate evaluation question-answering pairs for each facet based on the hallucination dataset as well as the factuality verification questions in Section 2.1. Then we also manually inspect their quality. One example of the evaluation QA pairs for each facet is shown in Figure 1 (More examples are in Appendix E). The specific prompt design for GPT-4o is shown in Appendix A.
@@ 2.3 Knowledge Editing Techniques
  • Fine-tuning is a simple and straightforward way to update the parametric knowledge of LLMs. We selected three variations FT-L (Meng et al., 2022), FT-M (Zhang et al., 2024b), and LoRA (Hu et al., 2022), which mitigate the catastrophic forgetting and overfiting issues of standard fine-tuning.
@@ 3.1 Facet 1: Efficacy
Figure 3 shows the Efficacy Score performance of post-edit LLMs in each domain and the whole HalluEditBench. Since we have ensured the LLMs’ pre-edit Efficacy Score is 0, Figure 3 can directly reflect the effectiveness of different knowledge editing techniques in correcting real-world hallucinations. Thus, we find that the effectiveness of some techniques can be far from what their performance on previous datasets suggests, implying the potential unreliability of performance on previous datasets. For example, as shown in Table 1, although FT-M achieves near $100\%$ performance in existing datasets such as $\text{WikiData}_{\text{recent}}$, ZsRE, and WikiBio, its overall Efficacy Scores on Llama2-7B and Mistral-v0.3-7B are only around $60\%$. There is a similar performance drop for MEMIT.
@@ 3.2 Facet 2: Generalization
Figure 4 shows both the pre-edit and post-edit Generalization Scores for different knowledge editing techniques on three LLMs. It is worth noting that the pre-edit performance is the same for different techniques and not zero, illustrating that the manifestation of hallucination actually depends on the design of question prompts. Given a group of diverse question prompts for the same knowledge triplet, LLMs may hallucinate some questions but answer others correctly.
@@
Comparing the ranking of Efficacy Scores in Figure 3 with Figure 4, we can explicitly see that higher Efficacy Scores do not also necessarily indicate higher Generalization Scores. Especially, although GRACE almost surpasses all the other editing techniques regarding Efficacy Scores, it largely degrades the Generalization Scores compared to pre-edit performance. In addition, all editing methods except ICE only marginally improve or even hurt Generalization Scores.
Figure 4: Generalization Scores of Knowledge Editing Methods. Generalization Scores (%) are measured by accuracy on five types of Generalization Evaluation Questions including Rephrased Questions (“rephrase”), Yes-or-No Questions with Yes or No as answers (“yes” or “no”), Multi-Choice Questions (“mc”), Reversed Questions (“reversed”). The “average” refers to averaged scores over five question types. The figure only shows the overall Generalization Scores for each type on the whole HalluEditBench. Generalization Scores for each domain are given in Appendix D.1.
Insight 2: (1) The manifestation of hallucination depends on question design; (2) Higher Efficacy Scores do not also necessarily indicate higher Generalization Scores; (3) All editing techniques except ICE only marginally improve or negatively impact the Generalization performance.
@@ 3.3 Facet 3: Portability
We surprisingly find that except that ICE may bring marginal improvement to the pre-edit performance, the other knowledge editing techniques even mostly underperform pre-edit Portability Scores, showing another type of negative effect of knowledge editing and LLMs cannot really reason with the edited knowledge in multi-hop questions regardless of knowledge editing methods. Comparing single-hop and multi-hop performance, we observe a sharp decrease for all the editing methods, which further underscores the challenges of answering multi-hop questions with edited knowledge.
Figure 5: Portability Scores of Knowledge Editing Methods. Portability Scores (%) are measured by the accuracy on Portability Evaluation Questions, which are Efficacy Evaluation Questions with $N$ hops ($N=1\sim 6$). The Portability Evaluation Questions are the same as Efficacy Evaluation Questions when $N$ is 1. The Portability Scores on two domains “human” and “places” are reported in the figure. The results for more domains are given in Appendix D.2. The “overall” refers to the Portability Score (%) on the whole HalluEditBench embracing 9 domains.
Insight 3: (1) LLMs may memorize answers rather than reason based on single-hop knowledge for multi-hop questions; (2) Editing methods except ICE mostly underperform pre-edit Portability Scores, implying LLMs cannot really reason with edited knowledge in multi-hop questions.
@@ 3.5 Facet 5: Robustness
First, based on overall Robustness Scores, we observe that LLMs themselves have a large impact on the robustness of edited knowledge. The same editing technique could show distinct trends as turns increase on different LLMs. For example, all editing methods have a sharp drop when turns go up on Llama2-7B, showing a low level of robustness. However, MEMIT, ROME, FT-M on Llama3-8B and MEMIT, ROME, FT-M, FT-L on Mistral-v0.3-7B maintain almost the same performance as turns increase, suggesting a relatively high level of robustness for the edited knowledge.
@@
Figure 7: Robustness Scores of Knowledge Editing Methods. Robustness Scores are calculated by the accuracy on Robustness Evaluation Questions with $M$ turns ($M=1\sim 10$). We regard Efficacy Scores as the Robustness Scores when $M$ is 0. The Robustness Scores on two domains “human” and “places” are reported in the figure. The results for more domains are given in Appendix D.3. The “overall” refers to the Robustness Score (%) on the whole HalluEditBench embracing 9 domains.
@@ 4 Related Work
Knowledge editing techniques have attracted increasing attention for their efficiency advantages in addressing obsolete or hallucinated information in LLMs (Wang et al., 2023b; Zhang et al., 2024b). In general, the existing editing techniques can be categorized into four types including Locate-then-edit (Meng et al., 2022; 2023), Fine-tuning based (Gangadhar & Stratos, 2024; Zhu et al., 2020; Wang et al., 2024a), In-Context Editing (Zheng et al., 2023; Shi et al., 2024; Fei et al., 2024), and Memory-based (Wang et al., 2024d; Hartvigsen et al., 2024; Mitchell et al., 2022; Yu et al., 2023). Recently, many benchmarks have been built to investigate the properties of knowledge editing from different perspectives (Rosati et al., 2024; Wu et al., 2023; Ge et al., 2024a; Ma et al., 2023; Wei et al., 2023; 2024a; Zhong et al., 2023; Lin et al., 2024; Huang et al., 2024; Liu et al., 2024b; Akyürek et al., 2023; Li et al., 2024a; d; 2023b; Gu et al., 2024; Powell et al., 2024). For example, Gu et al. (2024) proposed a benchmark to assess the side effect of 4 popular editing methods on 3 LLMs across 8 general capacity tasks. Rosati et al. (2024) built a new evaluation protocol to measure the efficacy and impact of knowledge editing in long-form generation. Wei et al. (2024a) introduced a multilingual knowledge editing benchmark embracing five languages. However, considering the fundamental motivation of applying knowledge editing to LLMs, which is to correct hallucinations, there is a pressing need to build a real-world hallucination dataset with rigorous verification and systematically analyze the performance of different editing methods. Thus, we proposed HalluEditBench to fill in the gap and provided new insights to facilitate the progress in the field of knowledge editing.
@@ Appendix A Reproducibility Statement
Given a subject and a relation, create 2-hop, 3-hop, 4-hop, 5-hop, and 6-hop questions, along with their correct answers.
Always use the provided subject and relation to create multi-hop questions, and avoid including any correct answers from other multi-hop questions.
Ensure the answers for multi-hop questions are correct, and do not use ‘N/A’ as answers. Output in JSON format. Below is an example:
@@
 “2hop_question”: “Who is the spouse of the Amazon founder?”,
 “2hop_answer”: “MacKenzie Scott”,
 “3hop_question”:  “Which university did the spouse of the Amazon founder attend for their undergraduate studies?”,
 “3hop_answer”:  “Princeton University”,
 “4hop_question”:  “In which city is the university that the spouse of the Amazon founder attended located?”,
 “4hop_answer”:  “Princeton”,
 “5hop_question”:  “In which state is the city located where the university that the spouse of the Amazon founder attended is situated?”,
 “5hop_answer”:  “New Jersey”,
 “6hop_question”:  “In which country is the state located where the city is situated that contains the university the spouse of the Amazon founder attended?”,
 “6hop_answer”:  “United States”,
@@ Appendix B Details of the Benchmarked Knowledge Editing Techniques
FT-M (Zhang et al., 2024b) In contrast to FT-L, which fine-tunes by maximizing the probability of all tokens in the target sequence based on the last token’s prediction, Fine-Tuning with Masking (FT-M) refines this approach to align more closely with the traditional fine-tuning objective. FT-M also targets the same FFN layer identified by causal tracing but employs a masked training strategy. Specifically, it uses cross-entropy loss on the target answer while masking out the original text, ensuring that the model is trained directly on the relevant target content. This approach mitigates potential deviations from the original fine-tuning objective and provides a more precise adjustment of the model’s weights with minimal disruption to unrelated model behavior.

### v2

# Can Knowledge Editing Really Correct Hallucinations?

@@ 1 Introduction
Large Language Models (LLMs) have shown superior performance in various tasks (Zhao et al., 2023). However, one critical weakness is that they may output hallucinations, referring to the non-factual information in generated content, for reasons such as the limit of models’ internal knowledge scope or fast-changing world facts (Zhang et al., 2023). Considering the high cost of retraining LLMs from scratch, knowledge editing has been designed as a new paradigm to correct erroneous or outdated factual knowledge in LLMs (Wang et al., 2023c).
@@
Although there are many existing question-answering datasets such as $\text{WikiData}_{\text{recent}}$ (Cohen et al., 2024), ZsRE (Yao et al., 2023), and WikiBio (Hartvigsen et al., 2024) widely used for knowledge editing evaluation, one common issue is that they do not verify whether LLMs, before applying knowledge editing, actually generate hallucinated answers to the evaluation questions. When such datasets are adopted to evaluate the performance of LLMs after being edited, it is hard to directly use the scores to judge the effectiveness of different knowledge editing techniques in correcting hallucinations, which is the motivation of applying knowledge editing to LLMs. To better illustrate this point, following the evaluation setting in  (Zhang et al., 2024e), we conducted a preliminary study to examine the pre-edit and post-edit performances of Llama2-7B on the aforementioned three evaluation datasets. As shown in Table 1, we can clearly observe that Llama2-7B achieves a relatively high performance, measured by the rate of answering the evaluation questions correctly (Accuracy (%)), even before applying knowledge editing techniques. Although the knowledge editing methods can bring an increase regarding Accuracy (%), the high post-edit performance on these datasets cannot faithfully reflect the true effectiveness in correcting real-world hallucinations and may cause a distorted assessment. Thus, the fundamental question remains insufficiently validated: Can knowledge editing really correct hallucinations in LLMs?
@@
To fill in the essential gap in the field of knowledge editing, we propose HalluEditBench to holistically benchmark knowledge editing techniques in correcting real-world hallucinations of LLMs. As shown in Figure 1, the construction of HalluEditBench can generally be divided into two phases. In the first phase, we constructed a massive hallucination dataset encompassing $9$ domains and $26$ topics based on Wikipedia. For each of Llama2-7B, Llama3-8B, and Mistral-v0.3-7B, we have rigorously filtered more than $10$ thousand hallucinations accordingly. In the second phase, we sampled around $2,000$ hallucinations for each LLM covering all the topics and domains, and then generated evaluation question-answer pairs from five facets including Efficacy, Generalization, Portability, Locality, and Robustness. Through extensive empirical investigation on performance of $7$ typical knowledge editing techniques, including FT-L (Meng et al., 2022), FT-M (Zhang et al., 2024e), MEMIT (Meng et al., 2023), ROME (Meng et al., 2022), LoRA (Hu et al., 2022), ICE (Zheng et al., 2023), and GRACE (Hartvigsen et al., 2024), regarding the aforementioned five dimensions, we have provided novel insights into their potentials and limitations. A summary of the insights is as follows:
@@
  • No editing methods can outperform others across five facets and the performance beyond Efficacy for all methods is generally unsatisfactory. Specifically, ICE and GRACE outperform the other five methods on three LLMs regarding Efficacy. All editing methods except ICE only slightly improve or negatively impact the Generalization performance. Editing techniques except ICE could even underperform pre-edit LLMs on Portability. FT-M and ICE surpass others on Locality performance. ICE has a poor Robustness performance compared to other methods.
@@ 2.2 Evaluation QA Pair Generation and Metrics
After constructing the hallucination dataset, we proposed to holistically assess the performance of knowledge editing methods in correcting hallucinations from five facets including Efficacy, Generalization, Portability, Locality, and Robustness. First, we leveraged GPT-4o to generate evaluation question-answering pairs for each facet based on the hallucination dataset as well as the factuality verification questions in Section 2.1. Then we also manually inspect their quality. One example of the evaluation QA pairs for each facet is shown in Figure 1 (More examples are in Appendix F). The specific prompt design for GPT-4o is shown in Appendix A.
@@ 2.3 Knowledge Editing Techniques
  • Fine-tuning is a simple and straightforward way to update the parametric knowledge of LLMs. We selected three variations FT-L (Meng et al., 2022), FT-M (Zhang et al., 2024e), and LoRA (Hu et al., 2022), which mitigate the catastrophic forgetting and overfiting issues of standard fine-tuning.
@@ 3.1 Facet 1: Efficacy
Since we have ensured that LLMs generate hallucinated answers to the Efficacy Evaluation Questions before editing, the pre-edit Efficacy Score for all editing techniques is 0. Thus, Efficacy Scores in Figure 3 can directly reflect the effectiveness of different techniques in correcting real-world hallucinations. We find that the effectiveness of some techniques can be far from what their performance on previous datasets suggests, implying the potential unreliability of their previous assessment. For example, as shown in Table 1, although FT-M achieves near $100\%$ performance in existing datasets such as $\text{WikiData}_{\text{recent}}$, ZsRE, and WikiBio, its overall Efficacy Scores on Llama2-7B and Mistral-v0.3-7B are only around $60\%$. There is a similar performance drop for MEMIT.
@@ 3.2 Facet 2: Generalization
As shown in Figure 4, even though the pre-edit Efficacy Score performances for different editing techniques on three LLMs are ensured 0, it is worth noting that the pre-edit Generalization Score performance is not 0 regarding each question type, illustrating that the manifestation of hallucination actually depends on the design of question prompts. Given a group of diverse question prompts for the same knowledge triplet, LLMs may hallucinate some questions but answer others correctly.
@@
Comparing the ranking of Efficacy Scores in Figure 3 with Figure 4, we can explicitly see that higher Efficacy Scores do not also necessarily indicate higher Generalization Scores. Especially, although GRACE almost surpasses all the other editing techniques regarding Efficacy Scores, it largely degrades the Generalization Scores compared to pre-edit performance. In addition, all editing methods except ICE only slightly improve or even hurt Generalization Scores.
Figure 4: Generalization Scores of Knowledge Editing Methods. Generalization Scores (%) are measured by accuracy on five types of Generalization Evaluation Questions including Rephrased Questions (“rephrase”), Yes-or-No Questions with Yes or No as answers (“yes” or “no”), Multi-Choice Questions (“mc”), Reversed Questions (“reversed”). The “average” refers to averaged scores over five question types. The figure only shows the overall Generalization Scores for each type on the whole HalluEditBench. Generalization Scores for each domain are given in Appendix E.1.
Insight 2: (1) The manifestation of hallucination depends on question design; (2) Higher Efficacy Scores do not also necessarily indicate higher Generalization Scores; (3) All editing techniques except ICE only slightly improve or negatively impact the Generalization performance.
@@ 3.3 Facet 3: Portability
We surprisingly find that except that ICE may bring marginal improvement to the pre-edit performance, the other knowledge editing techniques even mostly underperform pre-edit Portability Scores, showing another type of negative effect of knowledge editing and LLMs may not really reason with the edited knowledge in multi-hop questions for most knowledge editing methods. Comparing single-hop and multi-hop performance, we observe a sharp decrease for all the editing methods, which further underscores the challenges of answering multi-hop questions with edited knowledge.
Figure 5: Portability Scores of Knowledge Editing Methods. Portability Scores (%) are measured by the accuracy on Portability Evaluation Questions, which are Efficacy Evaluation Questions with $N$ hops ($N=1\sim 6$). The Portability Evaluation Questions are the same as Efficacy Evaluation Questions when $N$ is 1. The Portability Scores on two domains “human” and “places” are reported in the figure. The results for more domains are given in Appendix E.2. The “overall” refers to the Portability Score (%) on the whole HalluEditBench embracing 9 domains.
Insight 3: (1) LLMs may memorize answers rather than reason based on single-hop knowledge for multi-hop questions; (2) Editing methods marginally improve or degrade pre-edit Portability Scores, implying LLMs may not really reason with edited knowledge in multi-hop questions.
@@ 3.5 Facet 5: Robustness
First, based on overall Robustness Scores, we observe that LLMs themselves have a large impact on the robustness of edited knowledge. The same editing method could show distinct trends as turns increase on different LLMs. For example, all editing methods have a sharp drop when turns go up on Llama2-7B, showing a low level of robustness. However, MEMIT, ROME on Llama3-8B and Mistral-v0.3-7B maintain almost the same and relatively high performance as turns increase, suggesting a comparatively high level of robustness for the edited knowledge.
@@
Figure 7: Robustness Scores of Knowledge Editing Methods. Robustness Scores are calculated by the accuracy on Robustness Evaluation Questions with $M$ turns ($M=1\sim 10$). We regard Efficacy Scores as the Robustness Scores when $M$ is 0. The Robustness Scores on two domains “human” and “places” are reported in the figure. The results for more domains are given in Appendix E.3. The “overall” refers to the Robustness Score (%) on the whole HalluEditBench embracing 9 domains.
@@ 4 Related Work
Knowledge editing techniques have attracted increasing attention for their efficiency advantages in addressing obsolete or hallucinated information in LLMs (Wang et al., 2023c; Zhang et al., 2024e). In general, the existing editing techniques can be categorized into four types including Locate-then-edit (Meng et al., 2022; 2023), Fine-tuning based (Gangadhar & Stratos, 2024; Zhu et al., 2020; Wang et al., 2024a), In-Context Editing (Zheng et al., 2023; Shi et al., 2024; Fei et al., 2024), and Memory-based (Wang et al., 2024d; Hartvigsen et al., 2024; Mitchell et al., 2022; Yu et al., 2023). Recently, many benchmarks have been built to investigate the properties of knowledge editing from different perspectives (Rosati et al., 2024; Wu et al., 2023; Ge et al., 2024a; Ma et al., 2023; Wei et al., 2023; 2024a; Zhong et al., 2023; Lin et al., 2024; Huang et al., 2024c; Liu et al., 2024c; Akyürek et al., 2023; Li et al., 2024a; d; 2023b; Gu et al., 2024; Powell et al., 2024). For example, Gu et al. (2024) proposed a benchmark to assess the side effect of 4 popular editing methods on 3 LLMs across 8 general capacity tasks. Rosati et al. (2024) built a new evaluation protocol to measure the efficacy and impact of knowledge editing in long-form generation. Wei et al. (2024a) introduced a multilingual knowledge editing benchmark embracing five languages. However, considering the fundamental motivation of applying knowledge editing to LLMs, which is to correct hallucinations, there is a pressing need to build a real-world hallucination dataset with rigorous verification and systematically analyze the performance of different editing methods. Thus, we proposed HalluEditBench to fill in the gap and provided new insights to facilitate the progress in the field of knowledge editing.
@@ Appendix A Reproducibility Statement
Given a subject, a relation, a 1-hop question, and its answer, create 2-hop, 3-hop, 4-hop, 5-hop, and 6-hop questions, along with their correct answers.
Always use the provided subject and relation to create multi-hop questions and include the preceding question in the subsequent question (for example, include the 2-hop question in 3-hop question, include the 3-hop question in 4-hop question).
DO NOT include the correct answer to any previous multi-hop question in subsequent ones (for example, do not include the correct answer to the 2-hop question in the 3-hop or 4-hop questions).
Ensure that the answers for all multi-hop questions are accurate, and do not use ’N/A’ as an answer.
You must include the given subject and relation in all of the 2-hop, 3-hop, 4-hop, 5-hop, and 6-hop questions. Output in JSON format. An example is provided below:
@@
1hop_question: Who is the Amazon founder? 1hop_answer: Jeff Bezos
@@
 “2hop_question”: “Who is the spouse of the Amazon founder?”,  “2hop_answer”: “MacKenzie Scott”,
 “3hop_question”:  “Which university did the spouse of the Amazon founder attend for their undergraduate studies?”,  “3hop_answer”:  “Princeton University”,
 “4hop_question”:  “In which city is the university that the spouse of the Amazon founder attended located?”,  “4hop_answer”:  “Princeton”,
 “5hop_question”:  “In which state is the city located where the university that the spouse of the Amazon founder attended is situated?”,  “5hop_answer”:  “New Jersey”,
 “6hop_question”:  “In which country is the state located where the city is situated that contains the university the spouse of the Amazon founder attended?”,  “6hop_answer”:  “United States”,
@@ Appendix B Details of the Benchmarked Knowledge Editing Techniques
FT-M (Zhang et al., 2024e) In contrast to FT-L, which fine-tunes by maximizing the probability of all tokens in the target sequence based on the last token’s prediction, Fine-Tuning with Masking (FT-M) refines this approach to align more closely with the traditional fine-tuning objective. FT-M also targets the same FFN layer identified by causal tracing but employs a masked training strategy. Specifically, it uses cross-entropy loss on the target answer while masking out the original text, ensuring that the model is trained directly on the relevant target content. This approach mitigates potential deviations from the original fine-tuning objective and provides a more precise adjustment of the model’s weights with minimal disruption to unrelated model behavior.